OpenStudy (anonymous):

Determine the domain of the function. f(x) is equal to the square root of two minus x. x ≤ 2 All real numbers x > 2 All real numbers except 2

OpenStudy (anonymous):

\[f(x)=\sqrt{2-x}\]

jimthompson5910 (jim_thompson5910):

Hint: You cannot take the square root of a negative number (and get some real number)

jimthompson5910 (jim_thompson5910):

So the radicand (the stuff inside the square root) can't be negative

OpenStudy (anonymous):

can you please teach me how to do this step by step PLEASE

jimthompson5910 (jim_thompson5910):

Similar example: Find the domain of \[\Large g(x) = \sqrt{x+3}\] You can't take the square root of a negative number, so the radicand must be either 0 or positive. So... \[\Large x+3 \ge 0\] \[\Large x+3-3 \ge 0-3\] \[\Large x \ge -3\] is the domain

OpenStudy (anonymous):

but its positive ??

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean

OpenStudy (anonymous):

you said you can't take the square root of a negative number

jimthompson5910 (jim_thompson5910):

yes exactly, so whatever number that is in the square root is either a) zero or b) positive

OpenStudy (anonymous):

if your finding the domain of a square root your rules are that the number cant be negative. if you were taking the domain of a fraction the rules would be you cant have a zero on the denominator

OpenStudy (anonymous):

is the answer x ≤ 2

ganeshie8 (ganeshie8):

correct! good work !!

OpenStudy (anonymous):

i missed you @ganeshie8